A R-L circuit consists of one resistance 7.3 Ohm and one unknown Inductance. A voltage v=11 sin(319 t) is applied to this circuit and measured current is 9.19 A. Find out the unknown Inductance in mH unit.
Answer #1

from ohms law

v = i*Z

Z = sqrt(R^2+W^L^2)

V^2 = I^2(R^2+W^2L^2)

R^2+W^2L^2 = 121/9.19^2 = 1.43

W^2L^2 = 51.86

L^2 = 5.096*10^-4

L = 22.57 mH